• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. Cisbani

    Log entry time 15:56:14 on June19,2005

    Entry number 146620

    keyword=Shift Summary (Day 06/19)

    Shift Summary

    08:00-08:50 Trouble rebooting adaql2 (see Pete e-log 146570)

    08:55-09:30 Run 4701 (H2O flux=749.8)
    09:31-10:07 Run 4702 (H2O flux=750.1) no beam at the beginning for a couple of minutes
    10:08-10:36 Run 4703 (H2O flux=743.2)
    10:33-11:17 Run 4704 (H2O flux=749.0)
    11:17-11:50 Run 4705 (H2O flux=748.3)
    11:50-12:30 Run 4706 (H2O flux=746.5) Right-Septum tripped (around 12:00!), run stopped at 130k events

    12:31 Right Septum reset, when ramped-up tripped again (Upper coil over temperature alarm)

    12:50 PC called, he notice that the upper coil temperature is above 7.9 (upper limit). Suggest to switch off the beam in Hall A in order to speed up the cool down of the upper coil.

    13:10-13:20 Run 4707 RICH pedestal run, beam off

    13:22 Beam back at 55 uA

    13:27-13:30 Run 4708 (H2O flux=744.2); yunk run (forgot to re-configure RICHNF1)

    13:30-14:03 Run 4709 (H2O flux=751.3)

    14:03-14:25 Run 4710 (H2O flux=750.6)

    14:25 Problem with vacuum in the beam line. MCC turned off the beam and made a controlled access to restart the pump and check the gauge
    They will call us when the beam is available.

    14:30 Noticed that Right Septum quenched about 10 minutes ago.
    Alarm reset, magnet ramped up

    14:45 Pete suggest to reduce the pump speed at 600 (tacheometer)

    14:55 Beam back at 55 uA

    15:05-15:46 Run 4711 (H2O flux=600.2)

    15:46-15:50 Run 4712 (H2O flux=607.6)

    15:50 MCC informed about a new short controlled access to reset the same pump of the previous vacuum problem